# Gradient Guidance for Diffusion Models: An Optimization Perspective

### v1

@@ 5.1 Gradient-Guided Generation with A Pre-trained Score
Assume access to a pre-trained score network $s_{\theta}$ and gradient information of the objective function $f$. Let us present our Algorithm 1 that adapts the pre-trained diffusion model and iteratively updates the gradient guidance (14). The gradient guidance is able to steer the backward sampling process towards generating new samples with higher values of $f$. See Figure 1 for illustration.
Algorithm 1 takes as input any pre-trained score function $s_{\theta}(x,t)$ and adapts the backward sampling process with gradient guidance. In each iteration, it evaluates $\nabla f(\cdot)$ at samples generated from the previous iteration (Line 5(i)), and then computes the gradient guidance $\texttt{G}_{loss}$ using the newly queried gradient (Line 5(ii)). Using the updated gradient guidance, the backward process then generate new samples with improved objective values (Module 1). At the end of iterations, the algorithm outputs an adapted version of the diffusion model, specified by $(s_{\theta},\texttt{G}_{K})$, which generates samples with near-optimal objective values.
@@
It is worth highlighting that Algorithm 1 works with any pre-trained score network $s_{\theta}(x_{t},t)$. It retains the original score network and only updates the guidance term. The gradient guidance changes the generation process by an additive term to the backward SDE, without having to re-train the score network. Therefore, the algorithm is computationally efficient and easy to implement. We experimented with Algorithm 1 using a pre-trained score network with about 15M parameters; see Section 7 for details. In our experiment, a single run of the backward sampling process (Module 1) takes 4.6s, and Algorithm 1 takes 76min overall. Thus it is a rather light-weighted algorithm to implement and run.
5.2 Gradient-Guided Diffusion Converges to Regularized Optima
We analyze the convergence properties of Algorithm 1 and show that in final iterations, generated samples center around a regularized solution of the optimization objective $f$. Our theorems allow the pre-training data to have arbitrary distribution.
@@ 5.2 Gradient-Guided Diffusion Converges to Regularized Optima
The objective $f:\mathbb{R}^{D}\to\mathbb{R}$ is concave and $L$-smooth with respect to the (semi-)norm $\left\lVert\cdot\right\lVert_{\bar{\Sigma}^{-1}}$, i.e., $\left\lVert\nabla f(x_{1})-\nabla f(x_{2})\right\lVert_{\bar{\Sigma}}\leq L\left\lVert x_{1}-x_{2}\right\lVert_{\bar{\Sigma}^{-1}}$ for any $x_{1},x_{2}$.
While Algorithm 1 works with any pre-trained score network, we study its optimization properties focusing on the class of linear score functions given by
@@
Here (15) is a general linear function class. For comparison, a recent related paper Marion et al. (2024) assumes a more restricted class with $C_{t}\equiv I$ for when studying parameter optimization in diffusion models. With a linear score function (15), pre-training a diffusion model is essentially the same as using a Gaussian model to estimate the pre-training data distribution and then sampling from this estimated Gaussian. In this case, the guidance $\texttt{G}_{loss}$ is also linear in $x_{t}$, therefore the final output of the guided diffusion model also follows a Gaussian distribution; see (27) in Appendix C.
Recall we aim for an adapted diffusion model $(s_{\theta},\texttt{G}_{K})$ to generate samples with high values of $f$. Thus, we focus on the mean of the generated distribution (taking $T\to\infty$ in the backward sampling process of $(s_{\theta},\texttt{G}_{K})$), denoted by $\mu_{K}$, and establish its optimization guarantee.
Theorem 2 (Convergence to Regularized Maxima in Mean).
Proof Sketch
We observe that (18) resembles a gradient ascent update from $\mu_{k}\approx\bar{z}_{k}$ to $\mu_{k+1}$ corresponding to a regularzed optimization problem (17). In this regularized objective, the original objective $f(x)$ incorporates an additional proximal term with $\lambda:=1/\eta_{k}$. Therefore we can analyze the convergence of $\mu_{k}$ by following the classical argument for gradient optimization. The full proof is provided in Appendix C.1
Remarks.
This view of regularized optimization gives the following insights on gradient-guided diffusion models:
(i) The regularization term $\frac{\lambda}{2}\left\lVert x-\bar{\mu}\right\lVert^{2}_{\bar{\Sigma}^{-1}}$ in (17) is centered at the data’s mean $\bar{\mu}$. It penalizes samples that are far away from the pre-training data. The norm $\left\lVert\cdot\right\lVert_{\bar{\Sigma}^{-1}}$ suggests the regularization is strong in the direction where the original data distribution has low variance. In other words, it reveals that the pre-trained score function acts as a form of “prior” in the guided generation process. This prior favors samples that are proximal to its pre-training data distribution, even when additional guidance are present.
(ii) The regularization term cannot be made arbitrarily small. In particular, our theorem requires that $\lambda\geq L$. This demonstrates a limit of adapting diffusion models with guidance. As long as the score function stays unchanged, one cannot extrapolate from the pre-training data unlimitedly by solely adding gradient guidance. As a consequence, we cannot simply add gradient guidance to a diffusion model in order to reach the global maxima of any objective function. If the goal is to reach global optima, one has to update the pre-trained score network and refine it with newly collected data, we explore this approach in Section 6.
(iii) The linear convergence rate (16) is determined jointly by the smoothness of the objective function and strength of the regularization. We also pay a linear factor in the dimension $D$. In the following subsection, we will show that the gradient guidance $\texttt{G}_{loss}$ can reduce the dimension dependence from $D$ to $d$ if the data admits a latent low-dimensional subspace.
5.3 Gradient Guidance for Optimization in Latent Spaces
Next we focus on data with latent subspace as in Assumption 1. In the next theorem, we show that the generated distribution of our adapted model would converge, in expectation, to the maxima of a regularized version of $f$ within the subspace $\hbox{Span}(A)$.
@@ 5.3 Gradient Guidance for Optimization in Latent Spaces
$x^{*}_{A,\lambda}=\mathop{\mathrm{argmax}}_{x\in\hbox{Span}(A)}~{}\left\{f(x)-\displaystyle\frac{\lambda}{2}\left\lVert x-\bar{\mu}\right\lVert^{2}_{{\bar{\Sigma}}^{-1}}\right\}.$ (19)
Recall that the gradient guidance $\texttt{G}_{loss}$ is faithful to the data’s latent subspace, as proved in Theorem 1. As a result, the gradient-guided backward process maintains this latent subspace structure in the generated output. Therefore, all the generated samples and optimization iterates of Algorithm 1 belong to the latent subspace $\operatorname{Span}(A)$. In other words, the entire optimization process happens in the latent low-dimensional subspace. This facilitates a coherent and more efficient exploration in the solution space. Comparing to Theorem 2, the optimization gap in Theorem 3 is substantially smaller, reduced from $\mathcal{O}\left(D\right)$ to $\mathcal{O}\left(d\right)$. It means that the optimization process leverages the latent subspace and converges much faster.
Finally we note that Theorem 3 only establishes convergence in mean. The final output distribution of Algorithm 1, with a linear score function, is a Gaussian distribution supported on the latent subspace; see Appendix C equation (27).
@@ 6 Gradient-Guided Diffusion with Adaptive Fine-Tuning for Global Optimization
In the previous section, we have seen that adding guidance to a pre-trained diffusion model cannot improve the objective function unlimitedly. The pre-trained score function would act as a form of prior to keep the generated output proximal to the original data’s distribution. This leads to a regularization term in the optimization formulation.
Further, we consider adaptively fine-tuning the pre-trained diffusion model for generating samples to attain the unregularized global optima. The idea is not only to update the guidance in the backward sampling process but also to use generated samples to fine-tune the pre-trained score network. Empirically, fine-tuning diffusion models utilizing self-generated samples has been explored by Black et al. (2023); Clark et al. (2023).
@@ 6.1 Adaptive Fine-Tuning Algorithm with Gradient Guidance
We propose an adaptive version of the gradient-guided diffusion, where both the gradient guidance and the score networks are iteratively updated utilizing self-generated samples. The full algorithm is given in Algorithm 2.
@@
where $\mathcal{D}_{0}:=\mathcal{D}$ is the pre-training data. For illustration of this algorithm, please see also Figure 1.
In practice, to update the score network incorporating newly generated data, one does not have to exactly solve (20) by re-training the full model from scratch. Instead, (20) can be viewed as a guideline that motivates more computationally efficient ways for updating the pre-trained score. It is a common practice to only fine-tune the weights of the old model by performing gradient descent over a few batches of newly generated data, which is similar to the spirit of (20).
In our experiment, we implemented Algorithm 2 using a pre-trained U-net score function with 15M parameters and tested its performance on synthetic objectives. We implement the finetuning step by making one single Adam step over the new data. In our experiment, the iterated finetuning process of Algorithm 2 takes 91min overall, only slightly longer than the 76min taken by Algorithm 1. For details on the experiment results, please see Section 7.
@@ 6.2 Guided Generation Finds Unregularized Global Optima
where $\hat{C_{t}}$ is set to stay the same as in the pre-trained score and only $b_{t}$ gets updated during iterative fine-tuning. Marion et al. (2024) studied a similar function class where $\hat{C}_{t}$ is freezed to be $\hat{C}_{t}\equiv I$.
@@
where $f^{*}_{A}=\max\{f(x)|x\in\hbox{Span}(A)\}.$
The proof idea is similar to the proof of Theorem 2. For simplicity, we analyze the case where only the most recent sample batch $\mathcal{D}_{k}$ is merged with $\mathcal{D}_{0}$ for finetuning the score function. More specifically, we let $w_{k,i}=0$ for $0<i<k$ and $w_{k,0}=1-w_{k,k}$. Similar to the proof of Theorem 2, we obtain a recurisve update rule given by
where $\bar{z}_{k}\approx\mu_{k}$ is the empirical mean of previous samples. This update rule also closely resembles the gradient ascent iteration for maximizing a regularized objective. A key difference here is that we can control the weights $w_{k,i}$ to reduce the impact of $\mathcal{D}_{0}$ and make the regularization term vanish to zero. Thus the mean $\mu_{k}$ eventually converges to the global maxima. For the detailed arguments and proof of convergence, we refer readers to Appendix C.3.
Theorem 4 illustrates the effect of finetuning a diffusion model using self-generated data. For comparison, Theorem 3 showed that without finetuning the diffusion model can only generate new samples proximal to the pre-training distribution. Now if we allow finetuning using self-generated samples, the diffusion model can iteratively refines itself and reaches global optima, while preserving the latent subspace structure in its generated output.
Now let us we take on an optimization view. The convergence rate suggested by Theorem 4 matches with that of standard convex optimization, in terms of their dependence on $K$ the number of gradient evaluations. Further, if we compare the guided diffusion model with a standard gradient solver, the optimality gap of our algorithm scales with the small intrinsic dimension $\mathcal{O}(d)$, while standard gradient ascent converges much more slowly due to the large ambient dimension $\mathcal{O}(D)$. This comparison highlights the merits of “generative optimization”. More specifically, diffusion models leverage pre-training data to learn their intrinsic characteristics. Therefore, when we add gradient guidance to the pre-trained score function and use it for generation, it means that we are solving an optimization problem in its own intrinsic low-dimensional space. This leads to substantially more efficient exploration and faster convergence. This theoretical insight explains the practical successes of guided diffusion models on complex optimization problems, such as video creation, image synthesis and protein AI, where traditional methods do not work at all.
7 Numerical Experiments
7.1 Experiment Setup
We set the data’s ambient dimension as $D=64$ and the linear subspace dimension as $d=16$. The linear subspace is represented by an orthogonal matrix $A\in\mathbb{R}^{D\times d}$. We randomly generate a matrix $A$ and fix it once generated. After that, we sample a data point $X$ by first randomly sampling a latent variable $U\sim\mathcal{N}(0,I_{d})$ and computing $X=AU$. We independently sample a total of $65536$ data points as our pre-training data set. The objective functions considered in our experiments are $f_{1}(x)=10-(\theta^{\top}x-3)^{2}$ and $f_{2}(x)=5-0.5\lVert x-b\rVert.$ Here, $\theta$ and $b$ are randomly generated and fixed afterward. Since our data assumes a low-dimensional subspace representation, it is convenient to decompose $\theta$ into $\theta_{\perp}=(I-AA^{\top})\theta$ and $\theta_{\parallel}=AA^{\top}\theta$, representing the off-support and on-support components. We refer to $\frac{\lVert\theta_{\perp}\rVert}{\lVert\theta_{\parallel}\rVert}$ as the off/on-support ratio. Analogously, for a generated sample, we can also define its off/on-support ratio. Clearly, a small off/on-support ratio indicates close vicinity to the subspace.
Score Network Pre-training
We utilize a version of the U-Net (Ronneberger et al., 2015), with 14.8M trainable parameters. Note that this is a complicated network going beyond the linear score function class considered in our theories. Following the implementation of Denoising Diffusion Probabilistic Models (DDPM, Ho et al. (2020)), we train the U-Net o estimate the score function $\nabla\log p_{t}$, via minimizing the score matching loss introduced in Eqn. (3). We discretize the backward process to have $200$ time steps as in Nichol and Dhariwal (2021), and the U-Net is trained using our generated data set for 20 epochs. We use Adam as the optimizer, set the batch size as $32$, and set the learning rate to be $10^{-4}$. After the pre-training phase, we confirmed that the data subspace structure is well learned, as the generated samples using the pre-trained diffusion model have an average off/on-support ratio of $0.039$.
Implementation of Algorithm 1
Implementation of Algorithm 2
Algorithm 2 differs from Algorithm 1 in that it allows additional fine-tuning of the pre-trained score network. We adopt a computationally lightweight fine-tuning strategy: We only perform one Adam optimization step using the re-weighted loss given by Eqn. (20) with a batch of $32$ generated samples. We set the learning rate as $10^{-6}$. This simple strategy already demonstrates good performances as shown in Figure 7. Other implementation details are kept the same as those of Algorithm 1.
We run all experiments using one NVIDIA A100 GPU. Module 1 takes 4.6 seconds to generate a sample. Algorithm 1 takes 76 minutes, and Algorithm 2 takes 91 minutes.
7.2 Results
We first demonstrate our gradient guidance $\texttt{G}_{loss}$ preserves the subspace structure learned from the pre-trained model. For comparison, we also tested the naive guidance G defined following Lemma 1 (with $\Sigma=I$). For a quick reference, we repeat the definition here:
Comparing G and $\texttt{G}_{loss}$ on Preserving Subspace Structure
(a) Algorithm 1
(b) Algorithm 2
(c) 300-350 round of (a)
(d) 1000-1200 round of (b)
Algorithm 1 Converges to Regularized Optima
(a) $\theta=A\beta^{*}$
(b) $\frac{\|{\theta_{\bot}}\|}{\|{\theta_{\|}}\|}=9$
(c) $b=4\cdot\textbf{1}_{D}$
(d) $b\sim\mathcal{N}(4\cdot\textbf{1}_{D},9\cdot I_{D})$
Algorithm 2 Converges to Global Optima
@@ 7.2 Results
Algorithm 2 converges to the maximal value of the objective function $f_{1}=10-(\theta^{\top}x-3)^{2}$ as shown in Figure 7(a). In Figure 7(b), we visualize the distribution of generated samples of Algorithm 1 (blue) and 2 (red), respectively, as the iteration evolves. We see that samples from Algorithm 1 mostly stay close to the pre-training data distribution (area described by the dotted contour). In constrast, samples of Algorithm 2 move outside the contour, as the diffusion model gets finetuned using self-generated data.
(a) Convergence of Algorithm 2
(b) Distribution of generated samples
@@ 8 Conclusion
In this paper, we investigate the role and design of gradient guidance for adapting and fine-tuning a pre-trained diffusion model from an optimization perspective. We propose a gradient guidance based on a lookahead loss, as well as two variants of diffusion-based generative optimization algorithm utilizing such guidance. We provide optimization guarantees for adapting/fine-tuning diffusion models towards maximizing any target concave differentiable reward function. Our analysis has also been extended to linear subspace data, where our gradient guidance and adaptive algorithms provably preserve and leverage the latent subspace, thus they achieve faster convergence to near-optimal solutions.
Appendix A Characterization for Output Distribution of Backward Process

### v2

@@ 5.1 Gradient-Guided Generation with A Pre-trained Score
Assuming access to a pre-trained score network $s_{\theta}$ and the gradient of the objective function $f$, we present Algorithm 1 to adapt the diffusion model and iteratively update the gradient guidance (11). See Figure 1 for illustration.
Alg. 1 takes any pre-trained score function $s_{\theta}$ as input. Each iteration evaluates $\nabla f(\cdot)$ at samples from the previous iteration (Line 5(i)), computes the gradient guidance $\texttt{G}_{loss}$ with the new gradient (Line 5(ii)), and generates new samples using the updated guidance (Module 1). The algorithm outputs an adapted diffusion model specified by $(s_{\theta},\texttt{G}_{K})$.
@@
5.2 Gradient-Guided Diffusion Converges to Regularized Optima in Latent Space
We analyze the convergence of Alg. 1 and show that in final iterations, generated samples center around a regularized solution of the optimization objective $f$ within the subspace $\operatorname{Span}(A)$. Our theorems allow the pre-training data to have arbitrary distribution.
@@ 5.2 Gradient-Guided Diffusion Converges to Regularized Optima in Latent Space
The objective $f:\mathbb{R}^{D}\to\mathbb{R}$ is concave and $L$-smooth w.r.t. the (semi-)norm $\left\lVert\cdot\right\lVert_{\bar{\Sigma}^{-1}}$, i.e., $\left\lVert\nabla f(x_{1})-\nabla f(x_{2})\right\lVert_{\bar{\Sigma}}\leq L\left\lVert x_{1}-x_{2}\right\lVert_{\bar{\Sigma}^{-1}}$ for any $x_{1},x_{2}$.
While Alg. 1 works with any pre-trained score network, we study its optimization properties focusing on the class of linear score functions given by
@@
Remark on the linear parametrization of score network (12):
Analyzing the output distribution of guided diffusion is challenging because the additional guidance term destroys the dynamics of reverse SDE. A linear score is a natural and reasonable choice for characterizing the output distribution, as it was also adopted by [41].
With a linear score function (12), pre-training a diffusion model is equivalent to using a Gaussian model to estimate and sample from the estimated distribution. Thus, the guidance $\texttt{G}_{loss}$ is also linear in $x_{t}$, and the final output follows a Gaussian distribution; see (25) in Appendix D. We focus on the mean, $\mu_{K}$, of the generated distribution from the backward sampling of $(s_{\theta},\texttt{G}_{K})$ (as $T\to\infty$), and establish its optimization guarantee.
@@
$x^{*}_{A,\lambda}=\mathop{\mathrm{argmax}}_{x\in\operatorname{Span}(A)}~{}\left\{f(x)-\displaystyle\frac{\lambda}{2}\left\lVert x-\bar{\mu}\right\lVert^{2}_{{\bar{\Sigma}}^{-1}}\right\}.$ (13)
Remarks.
(1) The regularization term $\frac{\lambda}{2}\left\lVert x-\bar{\mu}\right\lVert^{2}_{\bar{\Sigma}^{-1}}$ (23) centers the data’s mean $\bar{\mu}$ and is stronger in directions where the original data has low variance. Thus, the pre-trained score acts as a "prior" in the guided generation, favoring samples near the pre-training data, even with guidance.
(2) The regularization term cannot be arbitrarily small, as our theorem requires $\lambda\geq L$. Thus, only adding gradient guidance cannot achieve the global maxima. If the goal is global optima, the pre-trained score must be updated and refined with new data, as explored in Section 6.
(3) The convergence rate is linear in the latent dimension $d$, rather than data dimension $D$. Since $\texttt{G}_{loss}$ is faithful to the latent subspace (Theorem 2), the generated samples and optimization iterates of Alg. 1 remain within $\operatorname{Span}(A)$. This leverage of the latent structure results in faster convergence.
@@ 6 Gradient-Guided Diffusion with Adaptive Fine-Tuning for Global Optimization
In the previous section, we have seen that adding guidance to a pre-trained diffusion model can’t improve the objective function unlimitedly due to the pre-trained score function acting as a regularizer. We consider adaptively fine-tuning pre-trained diffusion models to attain global optima. Empirically, fine-tuning diffusion models using self-generated samples has been explored by [7, 17].
@@ 6.1 Adaptive Fine-Tuning Algorithm with Gradient Guidance
We propose an adaptive version of the gradient-guided diffusion, where both the guidance and the score are iteratively updated using self-generated samples. The full algorithm is given in Algorithm 2.
@@
where $\mathcal{D}_{0}:=\mathcal{D}$ is the pre-training data. For illustration, please see also Figure 1, and the practical implementation of Alg. 2 is in Appendix E.
@@ 6.2 Guided Generation Finds Unregularized Global Optima
where $\hat{C_{t}}$ is set to stay the same in pre-trained scores, with only $b_{t}$ updated during iterative fine-tuning.
@@
where $f^{*}_{A}=\max\{f(x)|x\in\operatorname{Span}(A)\}.$
Theorem 3 illustrates that fine-tuning a diffusion model with self-generated data can reach global optima while preserving the latent subspace structure. The convergence rate matches standard convex optimization in terms of gradient evaluations, $K$. Compared to standard gradient solvers, guided diffusion models leverage pre-training data to solve optimization problems in a low-dimensional space, preserving desired structures and enabling more efficient exploration and faster convergence.
7 Experiments
7.1 Simulation
We conduct numerical simulations of Algorithms 1 and 2, following the subspace setup described in Assumption 1. Specifically, we set $d=16$, $D=64$, The latent variable $u$ is drawn from $\mathcal{N}(0,I_{d})$ and used to construct $x=Au$, where $A$ is a randomly generated orthonormal matrix. We define the objective function $f(x)=10-(\theta^{\top}x-3)^{2}$. To approximate the score function, we employ a version of the U-Net [51] with 14.8M trainable parameters. More details including how to set up $\theta$ are provided in Section E.1.
Figure 4: Comparison between two types of gradient guidance G and $\texttt{G}_{loss}$ (left: Alg. 1; right: Alg. 2). The off/on support ratio of the generated samples is defined as $r_{\mathrm{off}}=\frac{\|{x_{\bot}}\|}{\|{x_{\|}}\|}$.
Preserving Subspace Structure.
We first demonstrate that $\texttt{G}_{loss}$ preserves the subspace structure learned from the pre-trained model. For comparison, we also tested the naive guidance $\texttt{G}(x_{t},t):=\beta(t)\left(y-g^{\top}\mathbb{E}[x_{0}|x_{t}]\right)g$ (more details in Section E.1.). Figure 4 (left) shows that $\texttt{G}_{loss}$ performs much better than the naive gradient G in preserving the linear subspace. Figure 4 (right) demonstrates that off-support errors increase with adaptive score fine-tuning (Alg. 2) due to distribution shift, with G resulting in more severe errors than $\texttt{G}_{loss}$.
Convergence Results.
@@ 7.1 Simulation
Figure 5 (a) and (b) show that Alg. 1 converges to a sub-optimal objective value, leaving a gap to the maximal value. This aligns with our theory that the pre-trained model acts as a regularizer in addition to the objective function. Figure 5 (c) shows that Alg 2 converges to the maximal value of the objective function. As illustrated by Figure 5 (d), samples from Alg. 1 mostly stay close to the pre-training data distribution (dotted contour area), whereas samples from Alg. 2 move outside the contour as the diffusion model is fine-tuned with self-generated data.
(a) Alg. 1: $\theta=A\beta^{*}$
(b) Alg. 1: $\frac{\|{\theta_{\bot}}\|}{\|{\theta_{\|}}\|}=9$
(c) Convergence of Alg. 2
(d) Visualizing sample generattion
Figure 5: Convergence of Algorithms 1 and 2. (a) and (b) are under different $\theta$ for the objective function. (d) visualizes the distribution of the generated samples of Alg. 2 (red) and Alg. 1 (blue) across the iterations.
7.2 Image Generation
We validate our theory in the image domain for Algorithm 1. We employ the StableDiffusion v1.5 model [50] as the pre-trained model. For the reward model, we follow the approach outlined by [68] to construct a synthetic model. This model is based on a ResNet-18 [29] architecture pre-trained on ImageNet [20], with the final prediction layer replaced by a randomly initialized linear layer that produces scalar outputs. For more experiment details, refer to Section E.2.
Figure 6: Reward increase and effect on images across iterations. Left: Reward increases and converges across iterations. Larger guidance strength $y$ (smaller regularizer strength) results in higher convergent reward value. Right: Images become more abstract, shifting from realistic to virtual backgrounds as reward increases.
Results.
By Algorithm 1, the reward increases and converges. Figure 6 (left) shows the reward changes with optimization iterations. The hyperparameter $y$ tunes the strength of guidance and is inversely related to the strength of the regularizer (theoretical implications in Appendix D). A larger guidance strength (smaller regularizer strength) leads to a higher convergent reward value. Figure 6 (right) illustrates the changes in generated images across iterations. As the reward increases, the images become increasingly abstract, shifting from photo-realistic with detailed backgrounds to more virtual, stylized ones.
@@ 8 Conclusion
In this paper, we focus on gradient guidance for adapting or fine-tuning pre-trained diffusion models from an optimization perspective. We investigate the look-ahead loss based gradient guidance and two variants of diffusion-based generative optimization algorithms utilizing it. We provide guarantees for adapting/fine-tuning diffusion models to maximize any target concave differentiable reward function. Our analysis extends to linear subspace data, where our gradient guidance and adaptive algorithms preserve and leverage the latent subspace, achieving faster convergence to near-optimal solutions.
Appendix A Additional Related Works
Direct Latent Optimization in Diffusion Models.
Besides guidance methods, an alternative training-free route by optimizing the initial value of reverse process [61, 5, 33, 58, 46]. These methods typically backpropagate the gradient of reward directly to the initial latent vector through an ODE solver, utilizing the chain rule. Thus at inference time, the reverse process is unchanged except for being fed with an optimized initialization, different from the guidance method we studied.
Sampling and Statistical Theory of Diffusion Model
In contrast to the fruitful empirical advances, the theory of diffusion models is still limited. To the best of our knowledge, a theoretical understanding of fine-tuning diffusion models is absent. Existing results mainly focus on the sampling ability and statistical properties of unconditional diffusion models. In particular, for sampling ability, a line of works shows that the distribution generated by a diffusion model is close to the data distribution, as long as the score function is accurately estimated [19, 2, 8, 36, 12, 37]. The accuracy of the estimated score function is measured in terms of an $L_{\infty}$ or $L_{2}$-norm distance. More recently, [14, 13, 6] develop refined and tighter analyses using Taylor expansions of the discretized backward process and localization method. It is worth mentioning that the analysis in [14, 13, 6] extends to broader sample generation processes such as deterministic ones based on probabilistic ODEs. Going beyond distributions in Euclidean spaces, [18] analyzes diffusion models for sampling distribution supported on a low-dimensional manifold. Moreover, [43] consider sampling from symmetric spiked models, and [23] study sampling from Gibbs distributions using diffusion processes.
Turning towards the statistical theory of diffusion models, [56] and [38] provide asymptotic analyses, assuming a parametric form of the score function. Unfortunately, asymptotic analysis does not lead to concrete sample complexities. Later, concurrent works, [45] and [10], establish sample complexity bounds of diffusion models for estimating nonparametric data distributions. In high dimensions, their results highlight a curse of dimensionality issue without further assumptions, which also appears in [64] considering kernel methods. More interestingly, these works demonstrate that diffusion models can circumvent the curse of dimensionality issue if the data has low-dimensional structures. In the same spirit, [42] investigate learning high-dimensional graphical models using diffusion models, without the curse of dimensionality. For conditional diffusion models, [68, 25] establish sample complexity bounds for learning generic conditional distributions. We refer readers to [11] for an overview of contemporary theoretical progress.
Appendix B Characterization for Output Distribution of Backward Process